# TRA: Better Length Generalisation with Threshold Relative Attention

### v3

@@ Abstract
Transformers struggle with length generalisation, displaying poor performance even on basic tasks. We test whether these limitations can be explained through two key failures of the self-attention mechanism. The first is the inability to fully remove irrelevant information. The second is tied to position, even if the dot product between a key and query is highly negative (i.e. an irrelevant key) learned positional biases may unintentionally up-weight such information - dangerous when distances become out of distribution. Put together, these two failure cases lead to compounding generalisation difficulties. We test whether they can be mitigated through the combination of a) selective sparsity - completely removing irrelevant keys from the attention softmax and b) contextualised relative distance - distance is only considered as between the query and the keys that matter. We show how refactoring the attention mechanism with these two mitigations in place can substantially improve generalisation capabilities of decoder only transformers. [^1] Full code available at: https://github.com/microsoft/tra. Core mechanism also in Appendix B.
@@ 2.2 Contextual Relative Distance
${M}_{i,j}$ =$=\begin{cases}1&\text{if }S^{\prime}_{i,j}>0\\ 0&\text{otherwise}\end{cases}$ (4)
We are now going to compute our contextualised distance matrix, but only consider the relative distance to keys which make it through the threshold, this give us the contextualised relative distance. To do this we first initialise distances $\bm{\hat{D}}$ as ${\bm{M}}$. Consider this example:
$\bm{\hat{D}}$ $=\begin{matrix}1&0&0&0\\ 1&0&0&0\\ 0&1&1&0\\ 1&0&1&0\\ \end{matrix}$ (5)
We next perform a right-to-left cumulative sum within each row giving us our updated distances [^2] In this example we zero all indices that don’t meet the threshold for clarity, but note that this is not strictly necessary as they are masked in the operation from (10). $\bm{\bar{D}}$:
$\bm{\bar{D}}$ $=\begin{matrix}1&0&0&0\\ 1&0&0&0\\ 0&2&1&0\\ 2&0&1&0\\ \end{matrix}$ (6)
2.3 Parameterising D
Now we have arrived at our contextualised relative distances, but we still need some way of turning them into an actual weight. In order to do so we utilise a parameterised forget gate. Lets denote the residual stream for position i as ${\bm{x}}_{i}$, with ${\bm{W}}_{f}\in\mathbb{R}^{E\times 1}$ as the forget projection, $b$ as a scalar bias, and $\sigma$ as the sigmoid non-linearity. Then the weight for each position is given as:
@@ 2.3 Parameterising D
As each $\delta_{i}$ is a sigmoided scalar value, the more times it is multiplied with itself the smaller it gets. The closer $\delta_{i}$ is to 1 the longer the decay takes, the closer to zero the shorter. This provides the model a temporal memory. Enabling it to forget irrelevant past information if necessary, or to treat all timesteps uniformly if temporal ordering is not a concern. We opt for a recency bias for several reasons. Firstly, it has been generally shown to be useful when modeling natural language (Shen et al., 2018; Press et al., 2021; Zhu et al., 2021; Clark et al., 2025). Secondly, it is unclear whether a token occupying the ith position provides useful information beyond causal ordering. The first occurrence can be identified by proximity to the start of sequence token, and middle positions are better differentiated via their unique semantic context as described in Ebrahimi et al. (2024). Finally, decaying memory is the mechanism empowering the resurgence in state space models (Peng et al., 2023; Orvieto et al., 2023; De et al., 2024), recent successful recursive networks (Opper & Siddharth, 2025), and has shown promise with transformers (Csordás et al., 2021; Lin et al., 2025).
@@ 2.4 The Final Attention Weights
$\bar{A}_{ij}$ =$=\begin{cases}-1e11&\text{if }M_{ij}=0\\ \hat{A}_{ij}&\text{otherwise}\end{cases}$ (10)
@@
This gives us the full TRA mechanism. We use $\log(\bm{D^{\prime}})$ for improved numerical stability and weight scale following Lin et al. (2025). Note keys which do not meet the threshold are completely removed from the final softmax and do not count towards distance. This means that TRA exhibits both selective sparsity and allows semantic content to determine, and consequently synergise with, positional bias.
@@ 3.2 Baselines and Experimental Setup
Forgetting Transformer (FoT): The recent Forgetting Transformer (FoT) (Lin et al., 2025) utilises a forget gate as its positional encoding. The crucial difference between it and TRA is that the forgetting transformer uses standard relative distance compared with TRA’s contextualised alternative. It therefore serves as the truest test as to whether the issues we hypothesise regarding position are valid.
Differential Transformer (Diff): Introduced by Ye et al. (2024) the differential transformer utilises a twin heads approach where an auxiliary attention heads looks to perform noise reduction by down-weighting attention to irrelevant keys. It tests to what extent simply the ability to remove noise from the attention dot product is beneficial - without any further considerations regarding position.
Hyper-parameters: For all models we use the mini configuration from Turc et al. (2019): four heads, four layers and a 256 dimensional embedding size. The MLP is set 2x embedding size for both the linear and gate hidden units. Our focus is on small models following prior work (Zhou et al., 2023). Furthermore, attention glitches have been shown to persist at scale (Liu et al., 2023a), and the true solution for these tasks should not require additional layers. We use a linear warm-up for the first 5% of steps coupled with cosine decay. Other task and model specific hyper-parameters can be found in Appendix A.
@@ 3.3 Results
Induct OOD (50-100) $13.42\pm 1.96$ $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $7.85\pm 11.87$ $45.36\pm 43.99$ $15.00\pm 2.83$ $4.43\pm 0.0$ $20.42\pm 2.74$ $\mathbf{100\pm 0.0}$
Induct OOD (100-200) $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $\mathbf{99.90\pm 0.0}$
Induct OOD (200-300) $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $\mathbf{99.33\pm 0.0}$
@@
Copy OOD (50-100) $12.43\pm 1.66$ $0.0\pm 0$ 0.0 $3.1\pm 2.5$ $4.44\pm 4.79$ $26.97\pm 6.11$ $\mathbf{97.65\pm 4.7}$ $1.61\pm 0.29$ $86.47\pm 21.17$ $\mathbf{100\pm 0.0}$
Copy OOD (100-200) $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $66.04\pm 40.49$ $0.0\pm 0$ 0.0 $40.89\pm 31.27$ $\mathbf{99.87\pm 0.14}$
Copy OOD (200-300) $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $3.54\pm 4.35$ $0.0\pm 0$ 0.0 $2.08\pm 2.14$ $\mathbf{98.16\pm 1.82}$
@@
TRA succeeds: it is able to robustly generalise to all tasks. Its performance on copy and induct demonstrate strong capability for memory and associative recall. Its ability to solve flip-flops shows that it can combine positional and semantic information in a complementary fashion. To the best of our knowledge it is the first transformer network to fully solve the flip-flop benchmark of Liu et al. (2023a). It achieves 100% accuracy across seeds, which till this point has been only been achievable by LSTM networks. However, both FoT and CoPE come very close to full generalisation. To further tease apart performance differences on this core capability we generate an additional synthetic test in the same spirit as the original benchmark.
@@ 3.4 Flip-Flops++
Task: To further test the differences between contextualised positions and the baseline approaches we create a further synthetic test which we refer to as Flip-Flops++. The task is a follows: the input consists of instruction followed by a sequence of random symbols combined with a trigger symbol. The set of instructions are as follows: return the symbol that occurs immediately after the first occurrence of the trigger, immediately after it, and immediately before or after the last occurence. For example, if we use a as the trigger symbol an example could consist of: (instruction: before first, sequence: bcxaklcaztyab, out: x). The training set contains examples from each of the instructions sequence lengths 0-50. The test set consists of of sequences of lengths 51-500. This is a more challenging scenario than the original benchmark as dependency differences can be starker than in the original sparse set, and the model is required to behave differently depending on the instruction. Moreover, because we include having to retrieve from the first occurrence among the instructions it lets us investigate whether techniques which employ a recency bias are limited to only being able to attend to the most recent relevant item. For training we employ the same hyperparameters as before.
@@
Results are shown in Table 2. As before most PE methods struggle to generalise, though RoPE performs quite strongly with first match, echoing the findings of Barbero et al. (2025). However, in this more challenging setting, only the contextualised schemes, CoPE and TRA, are able to robustly generalise. Though TRA performs better at last match while CoPE holds a slight edge at first match, due to their use of different positional schemes. For natural language it is likely that the former is preferable as it better reflects causal structure (i.e. it allows understanding of an entity’s current state). Notably FoT, which in the original Flip-Flop benchmark of Liu et al. (2023a) appeared to be a strong contender, suffers substantial performance degradation in this more challenging setting. This is because, as we hypothesised in the introduction, it is unable to synergise positional and semantic information and consequently fails to generalise to capability (3) from earlier. This shortcoming is clearly visible by contrasting the attention weights of FoT and TRA.
@@ 3.5 Why does contextualised relative distance matter?
In order to analyse the impact of contextualised relative distance we train two layer one head TRA and FoT models on the flip-flop language modeling task and compare the attention heatmaps for the final layer in Figure 1. TRA fully generalises in this limited setting (the only model to be able to do so out of all the baselines) while FoT fails. Figure 1(a) shows how TRA is able to combine both semantic content with relative positions. It first filters out irrelevant keys fully and is then able to apply a strong positional bias to the remainder that truly matter, allowing it to cleanly select the most recent write instruction. Its recency bias (1(a) centre) is applied in a shifted manner only to what is relevant, which allows it express a strong positional preference. FoT on the other hand (1(b)), has no such mechanism and as a result must diminish its recency bias to be able to focus on semantic content (contrast the strength of 1(b) centre with 1(a) centre). This fact, coupled with its inability to completely remove irrelevant tokens, means that it is unable to handle long distance dependencies and consequently cannot learn semantic and positional preference such that they complement each other. We use FoT here as a clean illustrative example, but note that the same failure mode must be true of all schemes that do not employ contextualised distance.
@@ 3.6 Language Modeling:
Figure 2: WikiText-103 Test Perplexity ($\downarrow$) on OOD sequences lengths. Models trained for 100k steps with a window size of 128, and evaluated with window size 4098. Results taken over four random seeds.
Hyper-parameters: For language modelling we increase model size to the medium configuration of Turc et al. (2019). This is eight layers, eight heads and a 512 dimensional embedding size. The MLP is set 2x embedding size for both the linear and gate hidden units as before. We use the GPT-2 tokenizer (Radford et al., 2019) which leads to a vocabulary size of roughly 52k. This leads to the total number of parameters for both TRA and the baselines to be circa 100 million. We train using a window size of 128 and a batch size of 64 for 100k steps. Scheduling regime remains the same as before. Our evaluation consist of two parts. In distribution: we measure perplexity on the test set using the training window of size of 128. Out of distribution: we increase the window size to 4096 and observe the extent to which perplexity changes with increased length.
Results: In-distribution results are shown in table 3. TRA remains competitive with the strongest baselines. This means that we can introduce selective sparsity without hurting learning. Notably the strongest performing model in-distribution is the differential transformer (Ye et al., 2024), which contains a more sophisticated mechanism for relevancy determination. Replacing the simple ReLU we use here with such a more sophisticated approach could prove a fruitful avenue for future work. The out of distribution case is shown in figure 2. Most PE methods very rapidly begin to decline in quality, with similar degradation patterns emerging compared to the synthetic settings. The three chief exceptions to this rule are TRA, CoPE and FoT. CoPE is able to maintain/slightly improve perplexity for up to roughly a 10x increase in context length, before performance begins to deteriorate. We believe this is because its use of fractional positions only temporarily alleviates generalisation difficulties, but ultimately leaves the core issue unaddressed. Furthermore, increasing the number of positions in CoPE may delay degradation further, but any increase comes with a heavy computational burden and consequently makes CoPE difficult to scale as the sole mechanism for representing position. On the other hand, both FoT and TRA are capable of strong length generalisation (up to 32x greater than training length). Moreover, they display a promising trend of improved perplexity with increased length which indicates that they are able to make use of increased context rather than simply being robust to it. Between the two, TRA demonstrates a consistent perplexity edge compared with FoT. We believe this is due to its improved capacity for handling out of distribution dependency lengths as demonstrated by the flip-flop language modeling results. In sum, we show that TRA can introduce selective sparsity to the attention weights without compromising performance, and that its improved generalisation capabilities in synthetic settings also appear to be mirrored in the more complex setting of natural language.

### v4

@@ Abstract
Transformers struggle with length generalisation, displaying poor performance even on basic tasks. We test whether these limitations can be explained through two key failures of the self-attention mechanism. The first is the inability to fully remove irrelevant information. The second is tied to position, even if the dot product between a key and query is highly negative (i.e. an irrelevant key) learned positional biases may unintentionally up-weight such information - dangerous when distances become out of distribution. Put together, these two failure cases lead to compounding generalisation difficulties. We test whether they can be mitigated through the combination of a) selective sparsity - completely removing irrelevant keys from the attention softmax and b) contextualised relative distance - distance is only considered as between the query and the keys that matter. We show how refactoring the attention mechanism with these two mitigations in place can substantially improve the generalisation capabilities of decoder only transformers. [^1] Pytorch code for core mechanism in Appendix A.
@@ 2.2 Contextual Relative Distance
${M}_{i,j}$ $=\begin{cases}1&\text{if }S^{\prime}_{i,j}>0\\ 0&\text{otherwise}\end{cases}$ (4)
We are now going to compute our contextualised distance matrix, but only consider the relative distance to keys which make it through the threshold, this give us the contextualised relative distance. To do this we consider the following example $\bm{M}$}, where again only keys which meet the threshold are assigned 1:
${\bm{M}}$ $=\begin{bmatrix}1&0&0&0\\ 1&0&0&0\\ 0&1&1&0\\ 1&0&1&0\\ \end{bmatrix}$ (5)
Our threshold mask $\bm{M}$} is a $m\times n$ matrix, which we are going to use to get our contextualised relative distance matrix $\bm{\bar{D}}$, by applying the cumsum operation in the right to left direction (i.e. producing relative positions given causal masking). More formally we can say that for row index i and column index j:
$\bar{D}_{ij}={\sum}_{k=j}^{n}M_{ik}$ (6)
For our example $\bm{M}$} from (5) this would yield the following output, noting that we omit the keys which do not meet the threshold in this example for clarity:
$\bm{\bar{D}}$ $=\begin{bmatrix}1&0&0&0\\ 1&0&0&0\\ 0&2&1&0\\ 2&0&1&0\\ \end{bmatrix}$ (7)
$\bm{\bar{D}}$ now contains distances which are informed by context. For example, for the query in position two it views the key in position one as the closest possible item rather than its own key, which has been deemed irrelevant through the threshold operation. Similarly, for the query at position 4 it will view the key at position three as the closest possible item and the key at position 1 as the next nearest, skipping over the irrelevant intervening keys. This enables operation (3) from the introduction i.e. considering ordering but only over the relevant items. A capability which more traditional PE methods such as RPE (Dai et al., 2019), absolute (Vaswani et al., 2017) and RoPE Su et al. (2021) are unable to accomodate because they treat position as a fixed property rather than as adaptive to the context dependent needs of the query.
2.3 Parameterising $\bar{\bm{D}}$
Now we have arrived at our contextualised relative distances, but we still need some way of turning them into an actual weight. In order to do so we utilise a parameterised forget gate. Lets denote the residual stream of size E for position i as ${\bm{x}}_{i}$, with ${\bm{W}}_{f}\in\mathbb{R}^{E\times 1}$ as the forget projection, $b$ as a scalar bias, and $\sigma$ as the sigmoid non-linearity. Then the weight for each position is given as:
@@ 2.3 Parameterising $\bar{\bm{D}}$
As each $\delta_{i}$ is a sigmoided scalar value, the more times it is multiplied with itself the smaller it gets. The closer $\delta_{i}$ is to 1 the longer the decay takes, the closer to zero the shorter. This provides the model a temporal memory. Enabling it to forget irrelevant past information if necessary, or to treat all timesteps uniformly if temporal ordering is not a concern. We opt for a recency bias for several reasons. Firstly, it has been generally shown to be useful when modeling natural language (Shen et al., 2018; Zhu et al., 2021; Clark et al., 2025). In particular ALiBi (Press et al., 2021) shows that a non-parametric recency bias significantly improves length generalisation in transformers. However, though their non-parametric version makes this constraint rigid. By choosing to parameterise it instead we allow the model to completely disregard position if needed, the benefits of which are demonstrated by Lin et al. (2025). Secondly, it is unclear whether a token occupying the ith position provides useful information beyond causal ordering. The first occurrence can be identified by proximity to the start of sequence token. While middle positions are better differentiated via their unique semantic context, a capability which has been identified and carefully studied in transformers by Ebrahimi et al. (2024). Finally, a temporally decaying memory is a powerful mechanism. Merrill et al. (2024) show that is better able to approximate sophisticated $\text{NC}^{1}$ circuits given finite depth, compared to regular transformers. Moreover, it is the mechanism which has empowered the resurgence in state space models (Peng et al., 2023; Orvieto et al., 2023; De et al., 2024), recent successful recursive networks (Opper & Siddharth, 2024a), and has shown considerable promise with transformers (Csordás et al., 2021; Lin et al., 2025).
@@ 2.4 The Final Attention Weights
$\bar{A}_{ij}$ $=\begin{cases}-1e11&\text{if }M_{ij}=0\\ \hat{A}_{ij}&\text{otherwise}\end{cases}$ (11)
@@
This gives us the full TRA mechanism. We use $\log(\bm{D^{\prime}})$ for improved numerical stability and weight scale following Lin et al. (2025). Note keys which do not meet the threshold are completely removed from the final softmax by setting them to a large negative value (-1e11), as well as not counting towards distance as dicussed previously. This means that TRA exhibits both selective sparsity and allows semantic content to determine, and consequently synergise with, positional bias.
@@ 3.2 Baselines and Experimental Setup
Forgetting Transformer (FoT): The recent Forgetting Transformer (FoT) (Lin et al., 2025) utilises a forget gate as its positional encoding. The crucial difference between it and TRA is that the forgetting transformer uses standard relative distance compared with TRA’s contextualised alternative. This means that it can either enforce a recency bias or switch off any positional information completely. Crucially, what it cannot do is synergise the two (capability (3) from the introduction, which is precisely what TRA’s contextualised relative distance enables. FoT therefore serves as the truest test as to whether the issues we hypothesise regarding position are valid.
Differential Transformer (Diff): Introduced by Ye et al. (2024) the differential transformer utilises a twin heads approach where an auxiliary attention head looks to perform noise reduction by down-weighting attention to irrelevant keys. It tests to what extent simply the ability to remove noise from the attention dot product is beneficial - without any further considerations regarding position.
Hyper-parameters: For all models we use the mini configuration from Turc et al. (2019): four heads, four layers and a 256 dimensional embedding size. The MLP is set 2x embedding size for both the linear and gate hidden units. Our focus is on small models following prior work (Zhou et al., 2023). Furthermore, attention glitches have been shown to persist at scale (Liu et al., 2023a), and the true solution for these tasks should not require additional layers. We use a linear warm-up for the first 5% of steps coupled with cosine decay. Other task and model specific hyper-parameters can be found in Appendix B.
@@ 3.3 Results
Induct OOD (50-100) $13.42\pm 1.96$ $\pm 0.0 0.0±0.0$$\pm 0.0$ $7.85\pm 11.87$ $45.36\pm 43.99$ $15.00\pm 2.83$ $4.43\pm 0.0$ $20.42\pm 2.74$ $\mathbf{100\pm 0.0}$
Induct OOD (100-200) $\pm 0.0 0.0±0.0$$\pm 0.0$ $\pm 0.0 0.0±0.0$$\pm 0.0$ $\pm 0.0 0.0±0.0$$\pm 0.0$ $\pm 0.0 0.0±0.0$$\pm 0.0$ $\mathbf{99.90\pm 0.0}$
Induct OOD (200-300) $\pm 0.0 0.0±0.0$$\pm 0.0$ $\pm 0.0 0.0±0.0$$\pm 0.0$ $\pm 0.0 0.0±0.0$$\pm 0.0$ $\pm 0.0 0.0±0.0$$\pm 0.0$ $\mathbf{99.33\pm 0.0}$
@@
Copy OOD (50-100) $12.43\pm 1.66$ $0.0\pm 0.0$ $3.1\pm 2.5$ $4.44\pm 4.79$ $26.97\pm 6.11$ $\mathbf{97.65\pm 4.7}$ $1.61\pm 0.29$ $86.47\pm 21.17$ $\mathbf{100\pm 0.0}$
Copy OOD (100-200) $\pm 0.0 0.0±0.0$$\pm 0.0$ $\pm 0.0 0.0±0.0$$\pm 0.0$ $0.0\pm 0.0$ $66.04\pm 40.49$ $0.0\pm 0.0$ $40.89\pm 31.27$ $\mathbf{99.87\pm 0.14}$
Copy OOD (200-300) $\pm 0.0 0.0±0.0$$\pm 0.0$ $\pm 0.0 0.0±0.0$$\pm 0.0$ $0.0\pm 0.0$ $3.54\pm 4.35$ $0.0\pm 0.0$ $2.08\pm 2.14$ $\mathbf{98.16\pm 1.82}$
Results are in Table 1. We draw the following conclusions:
@@
TRA succeeds: it is able to robustly generalise to all tasks. Its performance on copy and induct demonstrate strong capacity for memory and associative recall. Its ability to solve flip-flops shows that it can combine positional and semantic information in a complementary fashion. To the best of our knowledge it is the first transformer network to fully solve the flip-flop benchmark of Liu et al. (2023a). It achieves 100% accuracy across seeds, which till this point has been only been achievable by LSTM networks. However, both FoT and CoPE come very close to full generalisation. To further tease apart performance differences on this core capability we generate an additional synthetic test in the same spirit as the original benchmark.
@@ 3.4 Flip-Flops++
Task: To further test the differences between contextualised positions and the baseline approaches we create a further synthetic test which we refer to as Flip-Flops++. The task is a follows: the input consists of instruction followed by a sequence of random symbols combined with a trigger symbol. The set of instructions are as follows: return the symbol that occurs immediately after the first occurrence of the trigger, immediately after it, and immediately before or after the last occurence. For example, if we use a as the trigger symbol an example could consist of: (instruction: before first, sequence: bcxaklcaztyab, out: x). The training set contains examples from each of the instructions and sequence lengths of 0-50. The test set consists of of sequences of lengths 51-500. This is a more challenging scenario than the original benchmark as dependency differences can be starker than in the original sparse set, and the model is required to behave differently depending on the instruction. Moreover, because we include having to retrieve from the first occurrence among the instructions it lets us investigate whether techniques which employ a recency bias are limited to only being able to attend to the most recent relevant item. For training we employ the same hyperparameters as before.
@@
Results are shown in Table 2. As before most PE methods struggle to generalise, though RoPE performs quite strongly with first match, echoing the findings of Barbero et al. (2025). However, in this more challenging setting, only the contextualised schemes, CoPE and TRA, are able to robustly generalise. Though TRA performs better at last match while CoPE holds a slight edge at first match, due to their use of different positional schemes. For natural language it is likely that the former is preferable as it better reflects causal structure (i.e. it allows understanding of an entity’s current state). Notably FoT, which in the original Flip-Flop benchmark of Liu et al. (2023a) appeared to be a strong contender, suffers substantial performance degradation in this more challenging setting. This is because, as we hypothesised in the introduction, it is unable to synergise positional and semantic information and consequently fails to generalise to capability (3) from earlier.
@@ 3.5 Why does contextualised relative distance matter?
In order to analyse the impact of contextualised relative distance we train two layer one head TRA and FoT models on the flip-flop language modeling task and compare the attention heatmaps for the final layer in Figure 1. TRA fully generalises in this limited setting, the only model to be able to do so out of all the baselines, and consistent with the optimal solution demonstrated in theory by Liu et al. (2023a). On the other hand FoT fails, and the reason why is clearly visible by contrasting the attention weights of the two.
@@
Figure 1(a) shows how TRA is able to combine both semantic content with relative positions. It first filters out irrelevant keys fully and is then able to apply a strong positional bias to the remainder that truly matter, allowing it to cleanly select the most recent write instruction. Its recency bias (1(a) centre) is applied in a shifted manner only to what is relevant, which allows it express a strong positional preference. FoT on the other hand (1(b)), has no such mechanism and as a result must diminish its recency bias to be able to focus on semantic content (contrast the strength of 1(b) centre with 1(a) centre). This fact, coupled with its inability to completely remove irrelevant tokens, means that it is unable to handle long distance dependencies and consequently cannot learn semantic and positional preference such that they complement each other. We use FoT here as a clean illustrative example, but note that the same failure mode must be true of all schemes that do not employ contextualised distance.
@@ 3.6 Language Modeling:
Hyperparameters: For language modelling we increase model size to the medium configuration of Turc et al. (2019); Opper et al. (2023a). This is eight layers, eight heads and a 512 dimensional embedding size. The MLP is set 2x embedding size for both the linear and gate hidden units as before. We use the GPT-2 tokenizer (Radford et al., 2019) which leads to a vocabulary size of roughly 52k. Totaling circa 80 million parameters for both TRA and the baselines. We train using: window size 128, batch size 64, 100k steps. Our scheduling regime remains the same as before. Our evaluation consist of two parts. In distribution: we measure perplexity on the test set using the training window of size of 128. Out of distribution: we increase the window size to 4096 and observe the extent to which perplexity changes with increased length.
@@
Figure 2: WikiText-103 Test Perplexity ($\downarrow$) on OOD sequences lengths. Models trained for 100k steps with a window size of 128, and evaluated with window size up to 4098. Results taken over four random seeds.
Results: In-distribution results are shown in table 3. TRA remains competitive with the strongest baselines. This means that we can introduce selective sparsity without hurting learning. Notably the strongest performing model in-distribution is the differential transformer (Ye et al., 2024), which contains a more sophisticated mechanism for relevancy determination. Replacing the simple ReLU we use here with such a more sophisticated approach could prove a fruitful avenue for future work. The out of distribution case is shown in figure 2. Most PE methods very rapidly begin to decline in quality, with similar degradation patterns emerging compared to the synthetic settings. The three chief exceptions to this rule are TRA, CoPE and FoT. CoPE is able to maintain/slightly improve perplexity for up to roughly a 10x increase in context length, before performance begins to deteriorate. We believe this is because its use of fractional positions only temporarily alleviates generalisation difficulties, but ultimately leaves the core issue unaddressed. Furthermore, increasing the number of positions in CoPE may delay degradation further, but any increase comes with a heavy computational burden and consequently makes CoPE difficult to scale as the sole mechanism for representing position. On the other hand, both FoT and TRA are capable of strong length generalisation (up to 32x greater than training length). Moreover, they display a promising trend of improved perplexity with increased length which indicates that they are able to make use of increased context rather than simply being robust to it. Between the two, TRA demonstrates a consistent perplexity edge compared with FoT. We believe this is due to its improved capacity for handling out of distribution dependency lengths as demonstrated by the flip-flop language modeling results. In sum, we show that TRA can introduce selective sparsity to the attention weights without compromising performance, and that its improved generalisation capabilities in synthetic settings also appear to be mirrored in the more complex setting of natural language. Finally, to provide an intuition for how TRA prefers to use its attention and forget gate we provide an analysis the attention heatmaps learned by our language models in Appendix C.